# Flow: A Modular Approach to Automated Agentic Workflow Generation

### v1

@@ 4 EXPERIMENTS
In all experiments, we compare Flowto the exists multi-agent frameworks i.e. (1) AutoGen [23], (2) Camel [8], and (3) MetaGPT [7]. In our experiments, we use agents empowered by GPT-4o-mini and GPT-3.5 [12].
@@
We designed three diverse and engaging tasks to evaluate multi-agent collaboration frameworks: 1) website development, 2) LaTeX Beamer slide creation, and 3) interactive game development. The rationale behind selecting coding-based experiments is twofold. First, most multi-agent frameworks tend to favour coding and writing abilities, like MetaGPT [7]. Using non-coding tasks may introduce bias. Second, coding tasks effectively demonstrate the system’s ability to assign agents and manage task allocation. Development of a Gobang Game with Naive AI: This task requires creating a Gobang (Five in a Row) game with a user interface and a simple AI opponent. Players can choose between black or white stones, with the UI clearly indicating turns and announcing the winner or draw when the game ends. This task demonstrates the system’s ability to handle modular design and task parallelism, as it involves coordinating game logic, AI implementation, and user interface development simultaneously.
Machine Learning Course Lecture Slides: This task focuses on generating LaTeX slides covering reinforcement learning algorithms, including motivations, problem statements, intuitive solutions, and detailed mathematical equations. A specific page requirement is to test the system’s ability to follow instructions precisely. The task highlights the system’s parallel processing capabilities of simultaneous generation of content, formatting, and presentation structure. The structured format of LaTeX also tests how effectively the system manages modularity and concurrent tasks.
Development of a Comprehensive Website for ICLR 2025: This task involves building a professional website for the International Conference on Learning Representations, hypothetically scheduled for San Francisco from April 27 to May 1, 2025. The website must feature key elements such as a detailed conference schedule and venue information with an interactive map. This task assesses each system’s ability to manage parallel workflows and modular components, including user interface design, functionality, and adherence to design guidelines, showcasing how well the system handles task decomposition and execution.
@@ 4.1 Evaluations over Three Designed Tasks
To conduct both quantitative and qualitative evaluations, we employed two metrics: Success Rate and Human Rating. Success Rate: The Success Rate is a quantitative measure that ranges from 0 to 1. Assesses whether the multi-agent system successfully generates executable outputs that fully meet the task requirements. A higher score indicates a greater level of success in accurately fulfilling the task objectives. Different tasks may have different evaluation metrics. The description for each evaluation metric is defined in Appendix B.1, B.2 and B.3. Human Rating: Human ratings are used to evaluate the quality of the generated results in alignment with the task description. We gathered 50 participants with programming and machine learning backgrounds to rank the outcomes produced by different methods. the detailed description of how we take scores is shown in the Appendix A
@@
We here give a summary of the performance of different methods over three tasks from Table LABEL:tab:game, 3 and 3, comparing the overall score regarding the success rate and human rating. The overall score of Flow  and human rating over three tasks, are (100, 4) on game design, (100, 3.33) on LaTeX writing, and (80, 3.28) on website design. Therefore, the average performance of Flow is 93% success rate and 3.54 over 4 satisfaction. Similarly, we have the average performance of AutoGen as (66.7, 2.75), MetaGPT as (71, 1.60), and CAMEL as (48.67, 2.12). Overall, our method Flow  has finished tasks with the most satisfaction and the highest success rate. Information about Flow’s workflow on those task is in Appendix D
4.2 Result for Gobang Game
The experimental setup is thoroughly detailed in Appendix B.2 and the visualisation result is in Fig.1. As shown in Table LABEL:tab:game, our method gets $100$ for all the aspects regarding success rate as well as the highest satisfaction from humans. More explanations for each method are as follows:
AutoGen: With the five tests, one trail failed to generate a valid output. Of the four successful attempts, one contained a code error that hindered normal execution, while another exhibited a bug in the game interface. The remaining two tests were completed successfully, though the chess pieces were displayed as the text “black” and “white”.
MetaGPT: After running MetaGPT five times, all attempts were successful and intractable. However, in four cases, a Tic-Tac-Toe game was generated instead of Gobang; out of these, the left one were functional, allowing both the user and AI to make moves and correctly terminate.
CAMEL: In all five trials, CAMEL was only successful twice. In the other attempts, the generated Python code was not executable. In the two successful trials, CAMEL successfully implemented correct termination conditions but had no AI component and terminated message.
Flow: After five rounds of testing, our system consistently generated successful outputs without any errors. The game functioned as expected, allowing both the player and the naive AI to take turns seamlessly. The game also ended correctly when either the board was fully occupied or one side achieved victory. In the game interface, the chess pieces were represented by actual black and white pieces, rather than text labels.
Table 1: Comparison of different LLM-based Multi-Agent frameworks on Gobang Game
@@ 4.2 Result for Gobang Game
Table 2: Comparison of different LLM-based Multi-Agent frameworks on LaTeX Beamer writing
@@
Table 3: Comparison of different LLM-based multi-agent frameworks on Website Design.
@@ 4.3 Result for LaTeX Beamer Writing
Experimental results are presented in Table 3 with explanations as follows:
AutoGen: After five tests, AutoGen successfully generated outputs every time. However, one output failed to compile in LaTeX due to syntax errors, and in two instances, the outputs did not meet the required length. The remaining outputs met both the length and content requirements.
MetaGPT: In five trials, four of them successfully generated a valid LaTeX version, with the only error being related to writing Python code within the ’.tex’ file. In these four successful trials, all documents met the required content specifications, but the total page count fell short of the requirements of 30 pages or 20 pages.
CAMEL: Successfully generated five different ’.tex’ files that are valid and could be rendered into Beamer format. Each presentation contained the required information, including sections like motivation. However, none of them met the page count requirement of 30 pages or 20 pages.
Flow: After five tests, our system successfully generated outputs each time, and all outputs were able to compile in LaTeX. However, one output contained repetitive content. In the remaining valid outputs, the length of the Beamer presentations met the specified requirements, and all the content mentioned in the requirements was adequately covered.
@@ 4.4 Result For Website Design
Similar to the previous two, the detailed experiment set-up is in Appendix B.3. We here illustrate the results in Table 3 as follows:
AutoGen: AutoGen produced HTML and CSS files with key information displayed but lacks details. Each section of the website contains only one or two sentences, lacking interactive functionality and necessary elements such as maps or tables.
MetaGPT: MetaGPT managed to create complete HTML and CSS, meeting basic functionality requirements and showcasing its code generation capabilities. However, the outputs were overly simplistic, missing content and key functional modules like the required venue and map.
CAMEL: CAMEL’s Outputs were executable in four out of five runs, though they did not include all the necessary elements, achieving all basic functions only. The system limits the communication can be only between two agents regardless of task complexity hindering its ability to fully complete complex website development tasks that require multi-task collaboration. Notably, one run generated complete HTML code but omitted the CSS file.
Flow: Flow achieves an 80% success rate within 5 trials. One trial failed to generate an HTML website. Among the remaining four trials, each section of the website featured detailed introductions and necessary interactive functionalities. For example, the venue section included travel information and local transportation options like airport, and accurately presented the conference location on a map. The registration section was fully functional, with a complete table, input boxes, and a submission button.
(a) Conference website: no newly added subtask, only the workflow is updated.
(b) Gobang Game: bad subtasks exist, add two new subtasks for successfully completing this task.
Figure 2: Workflow and dynamic update in two cases.
@@ 5 Workflow Update
Fig. 2(a) demonstrates the update process of Flow in the conference website creation example. Upon completing the first subtask, the system identifies potential changes and redundancies, triggering a restructuring process to enhance efficiency. Once the task ”Define the website structure” is completed, the generated data, which includes HTML structures and elements is sufficient to proceed with the CSS creation. As a result, the workflow is updated to incorporate the development of CSS based on the completed ”Define the website structure” task.
Fig. 2(b) illustrates a result of our dynamic updating process, where the system, upon receiving information about completed tasks, decides to add a bridging task to handle gaps and ensure the workflow continues smoothly.
@@
To evaluate the effectiveness of our updating mechanism, we intentionally introduced random masking to certain task outputs, replacing them with ”none” before passing them to the next agent. We conducted five trials and recorded the success scores. Since other frameworks employ a sequential workflow, we limit the comparison to our own approach in this context.
We observed a significant difference in success rate between using dynamic updating and not, particularly in the Interactive Game section as shown in Table 4. The main issue arises when the previous agent fails to provide the necessary information, yet the second agent continues with its task, leading to a major disconnect in the code. This often results in Python being unable to compile due to missing or mismatched components. Similarly, in website design, the lack of required elements caused by this failure impacts the overall functionality and structure. During the execution of subtasks, errors may arise due to the limitations of the LLM-based agent or underperformance in certain tasks. Therefore, the ability of updating agentic workflow dynamically to address such issues is essential.
@@ 6 Conclusion
We present Flow, a novel LLM-empowered multi-agent system that can dynamically adapt to unforeseen challenges for general tasks executions. With dynamically update the agentic workflow by AOV graphs, our system has largely fulfilling the modularity requirements for completing complex tasks. We demonstrate our method through case studies on a series of experiments, ranging from website design, game development and LaTeX Beamer creation as well as testing its capability on solving general benchmark tasks. Through objective evaluation metrics and human feedback, we found that Flow improves execution efficiency, offers better error tolerance, and delivers overall stronger performance.
@@ Appendix A Human Evaluation Process
Sometimes, LLMs can correctly fulfill each requirement of a task, but the quality of completion may vary. In such cases, human evaluation is necessary to assess the quality of the output. For each task, the final output of each Multi-Agent framework was evaluated by 50 participants, who ranked the outputs from best to worst. Points were awarded based on the rankings, with 1st place receiving 4 points and 2nd place receiving 3 points and so on. The final result was determined by calculating the average score. The detail distribution is shown in Fig. 4
@@
Figure 3: Ranking Distribution for conference website design. Shows that our results are better in the task of conference website design.
@@
Figure 4: Ranking Distribution for Gobang game making. Shows that our results are better in the task of Gobang game making.
@@ B.1 Experiment setup: LaTeX Beamer Writing
The task involves generating a LaTeX Beamer presentation, which is a popular LaTeX class used for creating professional-quality slides with various templates and effects. In this experiment, the objective is to produce presentations with different configurations, assessing the system’s ability to follow instructions. The experiment includes the following configurations:
@@ 5. Config 5:
    A 30-slide presentation, including motivation, problem statement, intuitive solution, and detailed mathematical equations.s
The goal is to examine the system’s ability to follow specific instructions while generating over 20 and 30 slides in different scenarios.
This task is well-suited for evaluation because it requires not only text generation but also an understanding of formatting and presentation logic. It serves as a comprehensive test of multitasking and reasoning capabilities. The structured nature of LaTeX allows for a rigorous assessment of the agent’s ability to manage complex, multi-component tasks, thereby highlighting the strengths of our method.
Evaluation Metrics: The following metrics are used to assess the performance of the generated LaTeX Beamer presentations:
@@
    Compilable: Verifies whether the LaTeX code compiles into a valid Beamer presentation. A successful compilation is rewarded with a score of 1, otherwise 0.
@@
    Completeness: Ensures that the final Beamer presentation includes all required components: motivation, problem, intuitive solution, and equations. Missing any of these results in a score of 0.
@@
The final result is calculated as the average of these three scores and shown in percentage.
B.2 Experiment setup: Go-bang game
@@
Gobang, also called Five in a Row, is a strategy board game where two players take turns placing black and white pieces on a grid. The objective is to be the first to align five consecutive pieces in a horizontal, vertical, or diagonal line. This experiment assesses our system’s ability to efficiently develop the game by utilizing parallelism to divide the development process into smaller, manageable tasks, such as game logic, AI move generation, and user interface (UI) design. We apply the same approach, taking the average score from five trials.
@@
    Compilable: The code compiles without errors. Any error that causes a termination will result in a score of 0.
@@
    Interactable: Properly supports both user and AI moves. If both functions are achieved score 1 else 0
@@
Each of these metrics is scored as 0 or 1, and the final result is calculated as the average of these scores and turn into percentage. These metrics allow for a comprehensive assessment of the efficiency, accuracy, and adaptability of each framework in developing a functional Gobang game with AI capabilities.
@@ B.3 Experiment setup: Website Design
We tasked the systems with developing a comprehensive website for the ICLR conference to evaluate their ability to handle complex tasks that require both flexible task coordination and effective problem-solving. This task tested the systems’ ability to manage multiple interdependent steps, such as designing user interfaces, ensuring functionality, and adhering to specific design guidelines.
@@
    Sections: Ensures inclusion of all required sections, with a focus on the schedule and venue as prompt asked. Missing required part in prompt will result of 0 in score.
By presenting a real-world scenario involving intricate requirements, we were able to observe how well the systems could break down a large project into manageable components and coordinate efforts across different tasks.
B.4 How Different LLMs Affect Updates
To verify how our framework performs with different capabilities of LLMs, we tested it on three tasks we designed using both GPT-4o-mini and GPT-3.5-Turbo. In this experiment, each task was run five times on different models, and the average of the results was calculated as the final outcome. We recorded three metrics: average init task, average changed task, and average changed ratio.
Init task refers to the number of tasks that need to be executed within the workflow after selecting the optimal workflow but before execution begins.
Average changed task indicates the number of tasks in the original workflow that were updated after completing the workflow execution.
Average changed ratio is calculated by dividing the average changed task by the init task, providing a more intuitive reflection of the proportion of tasks that were updated.

### v2

@@ 4 EXPERIMENTS
In all experiments, we compare Flowto the existing multi-agent frameworks i.e. (1) AutoGen [26], (2) Camel [10], and (3) MetaGPT [8]. In our experiments, we use agents empowered by GPT-4o-mini and GPT-3.5-Turbo [14].
@@
We designed three diverse and engaging tasks to evaluate multi-agent collaboration frameworks: 1) website design, 2) LaTeX Beamer writing, and 3) gobang game development. The rationale for selecting coding-based experiments is two-fold. First, most multi-agent frameworks, such as MetaGPT [8], are optimized for coding and writing tasks. Using non-coding tasks could introduce bias. Second, coding tasks effectively showcase the ability of a framework to assign agents and manage task allocation.
Gobang Game Development: This task requires creating a gobang game with a user interface and a simple AI opponent. Players can choose between black or white stones, with the UI clearly indicating turns and announcing the winner or draw when the game ends. This task demonstrates the framework’s ability to handle modular design and task parallelism, as it involves coordinating game logic, AI implementation, and user interface development simultaneously.
LaTeX Beamer Writing: This task focuses on generating LaTeX slides that cover reinforcement learning algorithms, including motivations, problem statements, intuitive solutions, and detailed mathematical equations. A specific page requirement is to test the framework’s ability to follow instructions precisely. The task highlights the framework’s parallel processing capabilities of simultaneous generation of content, formatting, and presentation structure. The structured format of LaTeX also tests how effectively the framework manages modularity and concurrent tasks.
Website Design: This task involves building a professional website for the International Conference on Learning Representations, hypothetically scheduled for San Francisco from April 27 to May 1, 2025. The website must feature key elements such as a detailed conference schedule and venue information with an interactive map. This task assesses each framework’s ability to manage parallel workflows and modular components, including user interface design, functionality, and adherence to design guidelines, showcasing how well the framework handles task decomposition and execution.
@@ 4.1 Evaluations over Three Designed Tasks
To conduct both quantitative and qualitative evaluations, we employ two metrics: Success Rate and Human Rating. The success rate is a quantitative measure that ranges from 0 to 1. It assesses whether the multi-agent framework successfully generates executable outputs that fully meet the task requirements. A higher score indicates a greater level of success in accurately fulfilling the task objectives. Different tasks may have different evaluation metrics. The description for each evaluation metric is defined in Appendix B.1, B.2 and B.3. Human ratings are used to evaluate the quality of the generated results in alignment with the task description. We gathered 50 participants with programming and machine learning backgrounds to rank the outcomes produced by different methods. A detailed description of how we take scores is shown in Appendix A.
@@
We summarize the performance of different methods on three tasks from Table LABEL:tab:game, 3 and 3, comparing the overall score with respect to the success rate and human rating. For Flow, the overall score and human rating over three tasks are (100, 4) on game development, (100, 3.33) on LaTeX writing, and (80, 3.28) on website design. Thus, the average performance of Flow is a 93% success rate and 3.54 out of 4 in human satisfaction. Similarly, we have the average performance of AutoGen as (66.7, 2.63), MetaGPT as (71, 1.60), and CAMEL as (48.67, 2.12). Overall, our method Flow has completed tasks with the most satisfaction and the highest success rate. Information about Flow’s workflow on those tasks is in Appendix D.
4.2 Result for Gobang Game Development
The experimental setup is thoroughly detailed in Appendix B.2 and the visualization result is shown in Fig.1. As shown in Table LABEL:tab:game, Flow  achieves a 100% success rate across all aspects, as well as the highest human satisfaction. More explanations for each method are given below.
AutoGen: Of the five trials, one of the tests failed to generate a valid result. Of the four successful attempts, one contained a code error that hindered normal execution, while another exhibited a bug in the game interface. The remaining two trials were completed successfully, although the chess pieces were displayed as the text ’black’ and ’white’ instead of graphical representations.
MetaGPT: After five trials, all MetaGPT attempts were successful and intractable. However, in four trials, a Tic-Tac-Toe game was generated instead of Gobang; out of these, the left one was functional, allowing both the user and AI to make moves and correctly terminate.
CAMEL: In all five trials, CAMEL was only successful twice. In the other trials, the generated Python code was not executable. In the two successful trials, CAMEL successfully implemented the correct termination conditions but did not have an AI component and no termination message.
Flow: After running Flow five times, our framework consistently generated successful outputs without errors. The game functioned as expected, allowing both the player and the naive AI to take turns seamlessly. The game also ended correctly when either the board was fully occupied or one side achieved victory. In the game interface, actual black and white chess pieces were displayed rather than text labels, enhancing the user experience.
Table 1: Comparison of different multi-agent frameworks on Gobang Game Development
@@ 4.2 Result for Gobang Game Development
Table 2: Comparison of different multi-agent frameworks on LaTeX Beamer Writing
@@
Table 3: Comparison of different multi-agent frameworks on Website Design
@@
(a) Website Design: no newly added subtask, only the workflow is updated.
(b) Gobang Game Development: adding two new subtasks to replace bad ones.
Figure 2: Workflow and dynamic update in two experiments.
@@ 4.3 Result for LaTeX Beamer Writing
Experimental results are presented in Table 3 with the following explanations:
AutoGen: After five trials, AutoGen successfully generated the output each time. However, one output failed to compile in LaTeX due to syntax errors, and in two instances, the outputs did not meet the required length. The remaining outputs met both the length and content requirements.
MetaGPT: In five trials, four of them successfully generated a valid LaTeX version, with the only error being related to writing Python code within the ’.tex’ file. In these four successful trials, all documents met the required content specifications, but only one meet the requirement of either 30 or 20 pages.
CAMEL: CAMEL successfully generated five valid ’.tex’ files, all of which could be rendered into Beamer format. Each presentation contained the required information, including sections such as motivation. However, none met the required page count of either 30 or 20 pages.
Flow: After five tests, our framework successfully generated output each time, and all outputs were able to be compiled in LaTeX. However, one output contained some repetitive content. In the remaining valid outputs, the Beamer presentations met the specified length requirements and adequately covered all required content.
@@ 4.4 Result For Website Design
Similarly to the previous two, the detailed experiment setup is in Appendix B.3. Here we illustrate the results in Table 3 as follows:
AutoGen: In five trials, four of the AutoGen results successfully rendered into an HTML website. However, in one attempt, each section of the website contained only one or two sentences and lacked interactive features and essential elements like maps or tables.
MetaGPT: MetaGPT managed to create complete HTML and CSS, meeting basic functionality requirements and showcasing its code generation capabilities. However, the outputs were overly simplistic, missing content and key functional sections like the required venue and map.
CAMEL: CAMEL’s outputs were executable in four out of five runs, though they did not include all the necessary elements, achieving all basic functions only. CAMEL restricts communication to only two agents, regardless of task complexity, hindering its ability to fully complete complex website development tasks. One of the results generated complete HTML code but omitted the CSS file.
Flow: Flow achieved an 80% success rate across five trials. One trial failed to generate an HTML website. Among the four remaining trials, each section of the website featured detailed introductions and necessary interactive functionalities. For example, the venue section included travel information and local transportation options. The registration section was fully functional, with a complete table, input boxes, and a submit button.
@@ 5 Workflow Update
Fig. 2(a) demonstrates the update process of Flow in the conference website creation example. Upon completion of the first subtask, the system identifies potential changes and redundancies, triggering a restructuring process to improve efficiency. Once the subtask ”Define the website structure” is completed, the generated data, which includes HTML structures and elements, is sufficient to proceed with the CSS creation. As a result, the workflow is updated to incorporate the development of CSS based on the completed ”Define the website structure” subtask.
Fig. 2(b) illustrates a result of our dynamic updating process, where the system, upon receiving information about completed subtasks, decides to add a bridging subtask to handle gaps and ensure that the workflow continues smoothly.
@@
To evaluate the effectiveness of our update mechanism, we intentionally introduced random masking to certain subtasks’ output, replacing them with ”none” before passing them to the next agent. We conducted five trials and recorded the success scores. Since other frameworks employ a sequential workflow, we limit the comparison to our own approach in this context.
We observed a significant difference in the success rate between using dynamic update and not, particularly in the Interactive Game section as shown in Table 4. The main issue arises when the previous agent fails to provide the necessary information, yet the second agent continues with its subtask, leading to a major disconnect in the code. This often results in Python being unable to compile due to missing or mismatched components. Similarly, in website design, the lack of required elements caused by this failure impacts the overall functionality and structure. During the execution of subtasks, errors may arise due to the limitations of the LLM-based agent or underperformance in certain tasks. Therefore, the ability to dynamically update the agent workflow to address such issues is essential.
@@ 6 Conclusion
We present Flow, a novel LLM-based multi-agent framework that can dynamically adapt to unforeseen challenges for general task executions. By dynamically updating the agentic workflow using AOV graphs, our framework has largely fulfilled the modularity requirements to complete complex tasks. We demonstrate our method through case studies on a series of experiments, ranging from website design, game development, and LaTeX Beamer writing, as well as testing its capability to solve general benchmark tasks. Through objective evaluation metrics and human feedback, we found that Flow  improves execution efficiency, offers better error tolerance, and delivers overall stronger performance.
@@ Appendix A Human Evaluation Process
Sometimes, LLM can correctly fulfill each requirement of a task, but the quality of completion may vary. In such cases, human evaluation is necessary to assess the quality of the output. For each task, the final output of each multi-agent framework was evaluated by 50 participants, who ranked the outputs from best to worst. Points were awarded based on the rankings, with the 1st place receiving 4 points, the 2nd place receiving 3 points, etc. The final result was determined by calculating the average score. The detailed distribution is shown in Fig. 4.
@@
Figure 3: Ranking distribution for website design across different frameworks. The results indicate that our method (Flow) outperforms others by achieving the highest percentage of first-place rankings.
@@
Figure 4: Ranking distribution for gobang game development across different frameworks. The results indicate that our method (Flow) outperforms others by achieving the highest percentage of first-place rankings.
@@ B.1 Experiment setup: LaTeX Beamer Writing
The task involves generating a LaTeX Beamer presentation, which is a popular LaTeX class used to create professional-quality slides with various templates and effects. In this experiment, the objective is to produce presentations with different configurations, assessing the framework’s ability to follow instructions. The experiment includes the following configurations:
@@ 5. Config 5:
The goal is to examine the framework’s ability to follow specific instructions while generating over 20 and 30 slides in different scenarios.
This task is well-suited for evaluation because it requires not only text generation but also an understanding of formatting and presentation logic. It serves as a comprehensive test of multitasking and reasoning capabilities. The structured nature of LaTeX allows for a rigorous assessment of the agent’s ability to manage complex, multicomponent tasks.
Evaluation Metrics: The following metrics are used to assess the performance of the generated LaTeX Beamer writing:
@@
    Compilable: Verifies whether the generated LaTeX code can compiles into a valid Beamer presentation or not. A successful compilation is rewarded with a score of 1, otherwise 0.
@@
    Completeness: Ensures that the final Beamer presentation includes all required components like: motivation, problem, intuitive solution, and equations. Missing any of these results in a score of 0.
@@
The final result is calculated as the average of these three scores and is shown as percentage.
B.2 Experiment setup: Gobang Game Development
@@
Gobang, also called ”Five in a Row”, is a strategy board game where two players take turns placing black and white pieces on a grid. The objective is to be the first to align five consecutive pieces in a horizontal, vertical, or diagonal line. This experiment assesses our framework’s ability to efficiently develop the game by utilizing parallelism to divide the development process into smaller, manageable tasks, such as game logic, AI move generation, and user interface (UI) design. We apply the same approach, taking the average score from five trials.
@@
    Compilable: The code compiles without errors. Any error that causes termination will result in a score of 0.
@@
    Interactable: Properly supports both user and AI movements. If both functions are achieved, score 1 else 0.
@@
Each of these metrics is scored as 0 or 1, and the final result is calculated as the average of these scores and turned into a percentage. These metrics allow for a comprehensive assessment of the efficiency, accuracy, and adaptability of each framework in developing a functional Gobang game with AI capabilities.
@@ B.3 Experiment setup: Website Design
We tasked the frameworks with developing a comprehensive website for the ICLR conference to evaluate their ability to handle complex tasks that require both flexible task coordination and effective problem solving. This task tested the ability of the frameworks to manage multiple interdependent steps, such as designing user interfaces, ensuring functionality, and adhering to specific design guidelines.
@@
    Sections: Ensures inclusion of all required sections, with a focus on the schedule and venue as prompt asked. Missing the required part in the prompt will result in a score of 0 in score.
By presenting a real-world scenario involving intricate requirements, we were able to observe how well the frameworks could break down a large project into manageable components and coordinate efforts across different tasks.
B.4 How Different LLM Affect Updates
To verify how our framework performs with different capabilities of LLM, we test both GPT-4o-mini and GPT-3.5-Turbo on three tasks we designed. In this experiment, each task was run five times on different models, and the average of the results was calculated as the final outcome. We recorded three metrics: average init task, average changed task, and average changed ratio.
Init task refers to the number of subtasks that need to be executed within the workflow after selecting the optimal workflow but before the execution begins.
Average changed task indicates the number of subtasks in the original workflow that were updated after the execution of the workflow.
Average changed ratio is calculated by dividing the average changed task by the init task, providing a more intuitive reflection of the proportion of subtasks that were updated.